OpenStudy (anonymous):

What values of x and y satisfy x + 2y = 6 and 2x + 3y = 4 simultaneously?

OpenStudy (australopithecus):

Use system of equations

OpenStudy (texaschic101):

are you familiar with the substitution method ?

OpenStudy (texaschic101):

x + 2y = 6 ----> x = -2y + 6 now sub -2y + 6 in for x in the other equation 2x + 3y = 4 2(-2y + 6) + 3y = 4 -- distribute through the parenthesis -4y + 12 + 3y = 4 -- subtract 12 from both sides -4y + 3y = 4 - 12 -- combine like terms -y = - 8 -- multiply entire equation by -1 to make y positive y = 8 now sub 8 in for y in either of the equations to find x x + 2y = 6 x + 2(8) = 6 x + 16 = 6 x = 6 - 16 x = -10 check ... 2x + 3y = 4 2(-10) + 3(8) = 4 -20 + 24 = 4 4 = 4 (correct) x = -10 and y = 8

OpenStudy (texaschic101):

do you have any questions ? Any at all ?